You plan build a sunroom on to your house in 6 months. You will pay for the sunroom in 3 installments of $10,000 each. The first payment is seven months from now and the other two payments are eight and nine months from now. If your investments earn 5% APR, compounded monthly, how much must you invest today to cover the cost of the sunroom?
Answer #1

Amount payable (p)=   10000
First payment due month (n)=   7
Second Payment due month (n)=   8
Third Payment due month=   9
investment rate = 5% compounded monthly  
monthly rate (i)= 0.05/12=   0.004166666667
  
Today invested Amount = sum of Present value of all amounts  
present value formula = P/(1+i)^n  
10000/(1+0.00416666667)^7+ 10000/(1+0.00416666667)^8+ 10000/(1+0.00416666667)^9  
=$29018.65977  
  
So, amount invested today is 29018.66